# Transfer Learning for High-dimensional Quantile Regression with Distribution Shift

### v1

@@ 2.1 Methodology of Tackling Three Types of Distribution Shift
Parameter Shift. Naturally we should consider a source study as informative if its parameter contrast $\boldsymbol{\delta}^{(k)}$ is relatively small. We choose the $\ell_{1}$-sparsity to quantify the informative level, as in many practical applications, parameter shift typically occurs across multiple dimensions without the overall magnitude growing too fast. Formally, the screening criterion is to reject source studies with $\|\boldsymbol{\delta}^{(k)}\|_{1}$ larger than an oracle transferring level, which has been widely adopted in the previous transfer work (Li et al.,, 2022; Tian and Feng,, 2023).
Residual Shift. Let’s start from the simplest low-dimensional linear model to understand how to tackle the residual shift. Suppose the target and source task as
$\left\{\begin{array}[]{l}\text{target model: }~{}y_{i}^{(0)}=(\boldsymbol{x}_{i}^{(0)})^{\top}\boldsymbol{\beta^{*}}+\epsilon_{i}^{(0)},~{}i=1,\dots,n,~{}\epsilon_{i}^{(0)}\sim\mathcal{N}(0,\sigma_{0}^{2}),\\ \text{source model: }~{}y_{i}^{(1)}=(\boldsymbol{x}_{i}^{(1)})^{\top}\boldsymbol{\beta^{*}}+\epsilon_{i}^{(1)},~{}i=1,\dots,n,~{}\epsilon_{i}^{(1)}\sim\mathcal{N}(0,\sigma_{1}^{2}).\end{array}\right.$
@@
$\left(\widehat{\boldsymbol{\beta}}_{tar}-\boldsymbol{\beta^{*}}\right)\sim\mathcal{N}\left(0,\frac{\sigma_{0}^{2}}{n}\Sigma_{x}^{-1}\right),\quad\left(\widehat{\boldsymbol{\beta}}_{pool}-\boldsymbol{\beta^{*}}\right)\sim\mathcal{N}\left(0,\frac{\sigma_{0}^{2}+\sigma_{1}^{2}}{4n}\Sigma_{x}^{-1}\right).$ (2)
If $\sigma_{1}^{2}$ is much larger than $\sigma_{0}^{2}$, i.e., the source study is much more noisy, then the pooling estimator will be much less efficient compared to the target one. From this perspective, addressing the residual shift is actually to avoid the increasing noise caused by incorporating divergent source studies.
Moving to the low-dimensional quantile regression with the same setting, we utilize the traditional quantile loss $\rho_{\tau}(x):=x(\tau-\mathbb{I}(x\leq 0))$ on the target data and the pooling data respectively. Based on the large sample theory of $M$-estimators (Van der Vaart,, 2000) and necessary regular conditions, it is easy to verify that
[…]
where $f^{(0)}(\cdot),f^{(1)}(\cdot)$ are density functions of residuals $\epsilon_{i}^{(0)},\epsilon_{i}^{(1)}$ respectively. When $f_{1}(0)\leq(\sqrt{2}-1)f_{0}(0)$, $\widehat{\boldsymbol{\beta}}_{pool}$ is less efficient than $\widehat{\boldsymbol{\beta}}_{tar}$. It can be seen that the residual density values $f^{(0)}(0)$ and $f^{(1)}(0)$ serve a similar role for QR estimators as residual variances $\sigma_{0}^{2}$ and $\sigma_{1}^{2}$ do for OLS estimators, both indicating the quality of information available for estimation. In fact, this relationship was evident in our motivating example, where unreliable source studies had considerably lower density values around the $\tau$-th quantile compared to the target study. When a source study has a very low density near the $\tau$-th quantile, its QR estimator often performs poorly due to the noisy information, making it unlikely that transferring from such a study would be beneficial. With this in mind, we propose another screening criterion to discard source studies with residual densities below a specified oracle threshold, as they are unlikely to contribute valuable information beyond introducing noise. For simplicity, we denote $f^{(k)}_{i}:=f^{(k)}(0|\boldsymbol{x}_{i}^{(k)})$ as the conditional density of residual $\epsilon_{i}^{(k)}$ at zero, hereinafter referred to as the $\tau$-th density.
@@
Rejecting sources with lower $\tau$-th density $f^{(k)}_{i}$, i.e., leveraging sources with higher $\tau$-th density, can also be explained from the linear regression example (2). If a source study has a negligible error term, i.e., $\sigma_{1}^{2}=0$, then it means accurately identifying the target parameter, leading to a wonderful transferable study. That is exactly why we choose this measure in the transferable set to address the residual shift. Additionally, with balanced study sizes, minimizing the estimator’s mean square error (MSE) benefits from a smaller contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}$ to reduce bias, while a larger $\tau$-th density helps to control variance. We will illustrate this point through theoretical results later in this section.
Transferable Set. From the analysis above, we need to control the transferring level for both parameter contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}$ and $\tau$-th density $f^{(k)}(0|\boldsymbol{x}_{i}^{(k)})$, in order to achieve efficient knowledge transfer in the presence of parameter and residual shift. Hence we propose the oracle transferable set $\mathcal{C}_{h}$ as defined in the introduction:
@@
$\mathcal{C}_{h}=\left\{k:\quad 1\leq k\leq K,\quad\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1},\quad\frac{\mathbb{E}[f^{(0)}_{i}]}{\mathbb{E}[f^{(k)}_{i}]}\leq h_{2}\right\},$ (3)
where $h_{1},h_{2}$ are oracle transferring levels revealing the magnitude of parameter and residual shifts respectively. One level $h_{1}$ controls the parameter similarity, while the other level $h_{2}$ regulates the expectation of source $\tau$-th density $\mathbb{E}[f^{(k)}_{i}]$ compared to the target $\mathbb{E}[f^{(0)}_{i}]$, facilitating the exclusion of heterogeneous source studies. Both requirements in the new transferable set (3) are actually elucidated in the earlier text. To guarantee the improvement of transfer learning, we only leverage the information from source studies included in this set $\mathcal{C}_{h}$. Strictly speaking, both transferring levels should be related to the sample size, since the impact caused by source studies differs with varying sample sizes under the same degree of shift. However, the impact from sample sizes is too hard to derive, where each level’s optimal choice depends on all target and source studies’ characteristics, as shown in the theoretical analysis later. Thus for simplicity, we set two universal oracle levels in (3) with implicitly assuming balanced sample sizes.
@@
Covariate Shift. The negative effect of covariate shift is mainly brought by the divergence of covariance matrices $\boldsymbol{\Sigma}^{(k)}$ between target and source studies. As mentioned by He et al., (2024), in the two-step transfer approach with the transferable set $\mathcal{C}_{h}$, the pooling estimator in the first step comes with an asymptotic bias $\boldsymbol{\delta}_{Pool}:=(\sum_{k\in\mathcal{C}_{h}}\boldsymbol{\Sigma}^{(k)})^{-1}\sum_{k\in\mathcal{C}_{h}}\boldsymbol{\Sigma}^{(k)}\boldsymbol{\delta}^{(k)}$. This bias can make the contrast $\boldsymbol{\delta}^{(k)}$ amplified by the factor $C_{\boldsymbol{\Sigma}}:=1+|\mathcal{C}_{h}|\max_{k\in\mathcal{C}_{h}}\sup_{1\leq j\leq p}\|e_{j}^{\top}(\boldsymbol{\Sigma}^{(k)}-\boldsymbol{\Sigma}^{(0)})(\sum_{k\in\mathcal{C}_{h}}\boldsymbol{\Sigma}^{(k)})^{-1}\|_{1}$, which may diverge when $\boldsymbol{\Sigma}^{(k)}$s are dissimilar. To address this problem, here we generalize the constrained optimization algorithm proposed by Li et al., (2024) to the field of quantile regression, for jointly estimating the target parameter $\boldsymbol{\beta^{*}}$ and contrast vectors $\{\boldsymbol{\delta}^{(k)}\}_{k\in\mathcal{C}_{h}}$. Our transfer framework is formalized as follows:
@@
subject to $\text{subject to }\left\{\begin{array}[]{l}\|S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{k},\quad\text{for }k\in\{0\}\cup\mathcal{C}_{h},\\ \|\sum_{k\in\{0\}\cup\mathcal{C}_{h}}S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{\boldsymbol{\beta}},\end{array}\right.$
where $S_{n_{k}}(\boldsymbol{\beta}):=\partial L^{(k)}(\boldsymbol{\beta})/\partial\boldsymbol{\beta}$ is the subgradient of the quantile loss function $L^{(k)}(\boldsymbol{\beta}):=\sum_{i=1}^{n_{k}}\rho_{\tau}(y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}\boldsymbol{\beta})$ for each study $k\in\{0\}\cup\mathcal{C}_{h}$, $\lambda_{\boldsymbol{\beta}},\{\lambda_{k}\}_{k\in\{0\}\cup\mathcal{C}_{h}}$ are tuning parameters. It is important to note that the $\ell_{1}$-sparsity regularizer for each contrast $\boldsymbol{\delta}^{(k)}$ allows us to be free from the similarity restriction of covariance matrices, hence removing the $C_{\boldsymbol{\Sigma}}$ factor and achieving robustness under covariance heterogeneity (Li et al.,, 2024; He et al.,, 2024). We will further reveal this property in the subsequent theoretical analysis.
@@
Here we give some explanations for our transfer framework. The objective function in (4) encourages sparse solutions of $\widehat{\boldsymbol{\beta}}$ and $\{\widehat{\boldsymbol{\delta}}^{(k)}\}_{k\in\mathcal{C}_{h}}$. The constraint $\|S_{n_{0}}(\boldsymbol{\beta})\|_{\infty}\leq\lambda_{0}$ is inherited from the target model, imposing that $\boldsymbol{\beta}$ should be identified as the true parameter $\boldsymbol{\beta^{*}}$ in the target task. The constraint $\|S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{k}$ comes from the $k$-th source model, imposing that $\boldsymbol{\delta}^{(k)}$ should be identified as the true contrast $\boldsymbol{w}^{(k)}-\boldsymbol{\beta^{*}}$. The last constraint $\|\sum_{k\in\{0\}\cup\mathcal{C}_{h}}S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{\boldsymbol{\beta}}$ aggregates target study and studies in the oracle transferable set $\mathcal{C}_{h}$, revealing that our estimation leverages information from valuable sources. In contrast to the two-step transfer framework (Li et al.,, 2022; Tian and Feng,, 2023), our algorithm can simultaneously estimate $\boldsymbol{\beta^{*}}$ and $\boldsymbol{\delta}^{(k)}$, providing more information for the downstream inference task.
Remark 3.
Zhao et al., (2023) proposed a residual importance weighted transfer approach to tackle distribution shift for high-dimensional linear regression. Specifically, they set the oracle weights as $\omega_{k}=f^{(0)}(y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}\boldsymbol{\beta^{*}})~{}/~{}f^{(k)}(y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}(\boldsymbol{\beta^{*}}+\boldsymbol{\delta}^{(k)}))$, then conduct the weighted LASSO to correctly identify $\boldsymbol{\beta^{*}}$ from the pooled data. Such weighting construction may still lose efficacy due to residual shift. If the source is abnormal like the motivating example, then the corresponding weight will approach infinity. However, we have shown that this source study ought to be discarded. Therefore, their approach may not be applicable to scenarios involving residual shift in our context.
To solve the constrained $\ell_{1}$-minimization problem (4), we approximate (4) with a joint $\ell_{1}$-penalized quantile regression ($\ell_{1}$-QR) form, leveraging the equivalence between Lasso and Dantzig selectors Candes and Tao, (2007); Bickel et al., (2009):
This formulation is similar to the co-training step in He et al., (2024), with replacing $\boldsymbol{\delta}^{(k)}$ by $\boldsymbol{w}^{(k)}-\boldsymbol{\beta}$. As suggested by Li et al., (2024), the algorithm for solving (5) proceeds naturally using iterative alternating direction methods. At a high level, our method works as follows. We first obtain an initial estimator $\widehat{\boldsymbol{\beta}}_{0}$ by $\ell_{1}$-QR solely on the target data (or the estimator $\widehat{\boldsymbol{\beta}}_{t-1}$ from the last iteration). Then for each source $k\in\mathcal{C}_{h}$ we subtract $(\boldsymbol{x}_{i}^{(k)})^{\top}\widehat{\boldsymbol{\beta}}_{t-1}$ from the response $y_{i}^{(k)}$ and obtain the estimator $\widehat{\boldsymbol{\delta}}^{(k)}_{t}$ by $\ell_{1}$-QR on the refined source dataset. After that, we take $\{\widehat{\boldsymbol{\delta}}^{(k)}_{t}\}_{k\in\mathcal{C}_{h}}$ into the objective function (5), solve for the estimator $\widehat{\boldsymbol{\beta}}_{t}$, and begin the next round. The iteration stops when estimators converge, where the detailed procedure is summarized in Algorithm 1. We defer the practical choice of tuning parameters $\lambda_{\boldsymbol{\beta}},\{\lambda_{k}\}_{k\in\{0\}\cup\mathcal{C}_{h}}$ to the simulation section.
Input: Target data $\{((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}\}$, source data $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k\in\mathcal{C}_{h}}$, tuning parameters $\lambda_{\boldsymbol{\beta}},\{\lambda_{k}\}_{k\in\{0\}\cup\mathcal{C}_{h}}$, iteration step $T$.
for $t=1:T$ do
  1. 1.
    Compute $\widehat{\boldsymbol{\delta}}^{(k)}_{t}\leftarrow\ell_{1}\text{-QR }\text{ on }[((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}\widehat{\boldsymbol{\beta}}_{t-1})_{i=1}^{n_{k}},\lambda_{k}],k\in\mathcal{C}_{h}$;
  2. 2.
    Solve the pooled $\ell_{1}$-QR problem

### v2

@@ 2.1 Methodology of Tackling Three Types of Distribution Shift
Parameter Shift. Naturally we should consider a source study as informative if its parameter contrast $\boldsymbol{\delta}^{(k)}$ is relatively small. We choose the $\ell_{1}$-sparsity to quantify the informative level, as in many practical applications, parameter shift typically occurs across multiple dimensions without the overall magnitude growing too fast. Formally, the screening criterion is to reject source studies with $\|\boldsymbol{\delta}^{(k)}\|_{1}$ larger than an oracle transferring level, which has been widely adopted in the previous transfer work [li2022transfer, tian2022transfer].
Residual Shift. We start from the simplest low-dimensional linear regression to understand how to tackle the residual shift. Suppose the target and source model as
$\left\{\begin{array}[]{l}\text{target model: }~y_{i}^{(0)}=(\boldsymbol{x}_{i}^{(0)})^{\top}\boldsymbol{\beta^{*}}+\epsilon_{i}^{(0)},~i=1,\dots,n_{0},~\epsilon_{i}^{(0)}\sim\mathcal{N}(0,\sigma_{0}^{2}),\\ \text{source model: }~y_{i}^{(1)}=(\boldsymbol{x}_{i}^{(1)})^{\top}\boldsymbol{\beta^{*}}+\epsilon_{i}^{(1)},~i=1,\dots,n_{1},~\epsilon_{i}^{(1)}\sim\mathcal{N}(0,\sigma_{1}^{2}).\end{array}\right$.
@@
$\left(\widehat{\boldsymbol{\beta}}_{tar}-\boldsymbol{\beta^{*}}\right)\sim\mathcal{N}\left(0,\frac{\sigma_{0}^{2}}{n_{0}}\Sigma_{x}^{-1}\right),\qquad\left(\widehat{\boldsymbol{\beta}}_{pool}-\boldsymbol{\beta^{*}}\right)\sim\mathcal{N}\left(0,\frac{n_{0}\sigma_{0}^{2}+n_{1}\sigma_{1}^{2}}{(n_{0}+n_{1})^{2}}\Sigma_{x}^{-1}\right).$
If $\sigma_{1}^{2}$ is much larger than $\sigma_{0}^{2}$, i.e., the source study is much more noisy, then the pooling estimator will be much less efficient compared to the target one.
Moving to the quantile regression with traditional loss $\rho_{\tau}(x):=x(\tau-\mathbb{I}(x\leq 0))$, assume $\epsilon_{i}^{(0)}\perp\boldsymbol{x}_{i}^{(0)},\epsilon_{i}^{(1)}\perp\boldsymbol{x}_{i}^{(1)}$ and $\mathbb{P}(\epsilon_{i}^{(0)}\leq 0)=\mathbb{P}(\epsilon_{i}^{(1)}\leq 0)=\tau$, denote the density functions of $\epsilon_{i}^{(0)},\epsilon_{i}^{(1)}$ as $f_{0}(\cdot),f_{1}(\cdot)$, respectively. Under some regular conditions, it is easy to verify that
$\left(\widehat{\boldsymbol{\beta}}_{tar}-\boldsymbol{\beta^{*}}\right)\xrightarrow{d}\mathcal{N}\left(0,\frac{\tau(1-\tau)}{n_{0}(f^{(0)}(0))^{2}}\Sigma_{x}^{-1}\right),~\left(\widehat{\boldsymbol{\beta}}_{pool}-\boldsymbol{\beta^{*}}\right)\xrightarrow{d}\mathcal{N}\left(0,\frac{\tau(1-\tau)(n_{0}+n_{1})}{(n_{0}f^{(0)}(0)+n_{1}f^{(1)}(0))^{2}}\Sigma_{x}^{-1}\right).$
If $f^{(0)}(0)$ is much smaller than $f^{(1)}(0)$, for example, when $n_{1}f^{(1)}(0)\leq(\sqrt{1+n_{1}/n_{0}}-1)n_{0}f^{(0)}(0)$, $\widehat{\boldsymbol{\beta}}_{pool}$ will be less efficient than $\widehat{\boldsymbol{\beta}}_{tar}$. The residual density values $f^{(0)}(0)$ and $f^{(1)}(0)$ serve a similar role for QR estimators as residual variances $\sigma_{0}^{2}$ and $\sigma_{1}^{2}$ do for OLS estimators, both indicating the quality of information available for estimation. Also, information quality depends on the sample size, since given the density function, a larger sample size provides richer quantile information. Consequently, another screening criterion should exclude source studies with small sample sizes $n_{k}$ and low densities at the $\tau$-th quantile $f^{(k)}(0|\boldsymbol{x}_{i}^{(k)})$, as they are unlikely to contribute valuable information beyond introducing noise. For simplicity, we denote $f^{(k)}_{i}:=f^{(k)}(0|\boldsymbol{x}_{i}^{(k)})$ as the conditional density of residual $\epsilon_{i}^{(k)}$ at zero, hereinafter referred to as the $\tau$-th density.
@@
Rejecting sources with lower $\tau$-th density $f^{(k)}_{i}$, i.e., leveraging sources with higher $\tau$-th density, can also be explained as reducing external variance. If a source study has a negligible variance term, i.e., $\sigma_{1}^{2}=0$, then it means accurately identifying the target parameter, leading to a wonderful transferable study. From the perspective of minimizing the mean square error (MSE), the transfer estimator benefits from a smaller contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}$ to reduce bias, while a larger $\tau$-th density helps to control variance. We further extend this motivating case to the covariate shift setting, where similar insights hold; please see Section 8.1.1 of supplementary materials for details.
Building on the previous analysis, we propose the oracle transferable set $\mathcal{C}_{h}$ as:
@@
$\mathcal{C}_{h}=\left\{k:\quad 1\leq k\leq K,\quad\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1},\quad\frac{n_{0}\mathbb{E}[f^{(0)}_{i}]}{n_{k}\mathbb{E}[f^{(k)}_{i}]}\leq h_{2}\right\},$ (2)
where $h_{1}$ and $h_{2}$ are oracle transferability levels characterizing the magnitude of parameter and residual shift, respectively. For each source study, the term $\|\boldsymbol{\delta}^{(k)}\|_{1}$ reflects parameter similarity, while the term $n_{k}\mathbb{E}[f^{(k)}_{i}$] captures the information quality around the objective quantile. The multiplicative form of $n_{k}\mathbb{E}[f^{(k)}_{i}$] is motivated by the structure $\sum_{k\in{0}\cup\mathcal{C}_{h}}n_{k}\mathbb{E}[f_{i}^{(k)}]$ appearing in Theorem 1 in the next subsection. To guarantee the improvement of transfer learning, we only leverage the information from source studies included in this set $\mathcal{C}_{h}$. It is important to note that, unlike $h_{1}$, which is expected to converge to zero as in the previous literature [li2022transfer, tian2022transfer], $h_{2}$ depends on the sample sizes of the target and source studies, thus is not necessary to converge to zero or diverges to infinity.
@@
Covariate Shift. The negative effect of covariate shift is mainly brought by the divergence of covariance matrices $\boldsymbol{\Sigma}^{(k)}:=\mathbb{E}[\boldsymbol{x}_{i}^{(k)}(\boldsymbol{x}_{i}^{(k)})^{\top}]$ between target and source studies. As mentioned by [li2022transfer, he2024transfusion], in the two-step transfer approach with transferable set $\mathcal{C}_{h}$, the pooling estimator in the first step makes the contrast $\boldsymbol{\delta}^{(k)}$ amplified by the factor $C_{\boldsymbol{\Sigma}}:=1+|\mathcal{C}_{h}|\max_{k\in\mathcal{C}_{h}}\sup_{1\leq j\leq p}\|\boldsymbol{e}_{j}^{\top}(\boldsymbol{\Sigma}^{(k)}-\boldsymbol{\Sigma}^{(0)})(\sum_{k\in\mathcal{C}_{h}}\boldsymbol{\Sigma}^{(k)})^{-1}\|_{1}$, which may diverge when source covariance matrices $\boldsymbol{\Sigma}^{(k)}$ are dissimilar to the target $\boldsymbol{\Sigma}^{(0)}$. To address this problem, here we generalize the constrained optimization algorithm proposed by [li2023estimation] to the field of quantile regression, for jointly estimating the target parameter $\boldsymbol{\beta^{*}}$ and contrast vectors $\{\boldsymbol{\delta}^{(k)}\}_{k\in\mathcal{C}_{h}}$. Our transfer framework is formalized as:
@@
$\text{subject to }\left\{\begin{array}[]{l}\|S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{k},\qquad\text{for }k\in\{0\}\cup\mathcal{C}_{h},\\ \|\sum_{k\in\{0\}\cup\mathcal{C}_{h}}S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{\boldsymbol{\beta}},\end{array}\right$.
where $S_{n_{k}}(\boldsymbol{\beta}):=\partial L^{(k)}(\boldsymbol{\beta})/\partial\boldsymbol{\beta}$ is the subgradient of the quantile loss function $L^{(k)}(\boldsymbol{\beta}):=\sum_{i=1}^{n_{k}}\rho_{\tau}(y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}\boldsymbol{\beta})$, and $\lambda_{\boldsymbol{\beta}},\{\lambda_{k}\}_{k\in\{0\}\cup\mathcal{C}_{h}}$ are tuning parameters. To solve this constrained $\ell_{1}$-minimization problem (3), we can follow the iterative algorithm adopted in [li2023estimation], where the detailed procedure is summarized in Section 8.1.2 of supplementary materials. Notably, the $\ell_{1}$-sparsity regularizer for each contrast $\boldsymbol{\delta}^{(k)}$ allows us to be free from the similarity restriction of covariance matrices, hence removing the $C_{\boldsymbol{\Sigma}}$ factor and achieving robustness under covariance heterogeneity [li2023estimation, he2024transfusion].
@@
In our transfer framework (3), the objective function encourages sparse solutions of $\widehat{\boldsymbol{\beta}}$ and $\{\widehat{\boldsymbol{\delta}}^{(k)}\}_{k\in\mathcal{C}_{h}}$. The constraint $\|S_{n_{0}}(\boldsymbol{\beta})\|_{\infty}\leq\lambda_{0}$ is inherited from the target model, imposing that $\boldsymbol{\beta}$ should be identified as the true parameter $\boldsymbol{\beta^{*}}$ in the target task. The constraint $\|S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{k}$ comes from the $k$-th source model, imposing that $\boldsymbol{\delta}^{(k)}$ should be identified as the true contrast $\boldsymbol{w}^{(k)}-\boldsymbol{\beta^{*}}$. The last constraint $\|\sum_{k\in\{0\}\cup\mathcal{C}_{h}}S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{\boldsymbol{\beta}}$ aggregates target study and studies in the oracle transferable set $\mathcal{C}_{h}$, revealing that our estimation leverages information from valuable sources. In contrast to the two-step transfer framework [li2022transfer, tian2022transfer], our algorithm can simultaneously estimate $\boldsymbol{\beta^{*}}$ and $\boldsymbol{\delta}^{(k)}$, providing more information for the downstream inference task.